Light falling on metallic surface of work function 2eV is given by B=100[sin(5×10^15)t +cos(8×10^15)t]. Then maximum kinetic energy of the photoelectrons is______

Grade:12

1 Answers

Sriram
5 years ago
Omega=5×10^15 and B0=100 
As kmax=homega-2π/2π (by applying v=omega/2π)
On substituting the values we get 
3.3ev